Functions f , g : R → R are defined, respectively, by f (x) = x2 + 3x + 1, g (x) = 2x – 3,
find  g o f  

Grade:12

1 Answers

Roshan bist
3 years ago
Gof(X)=g(f(X)=g(x^2+3x+1)=2(x^2+3x+1)-3=4x^2+6x+2-3=4x2+6x-1 is the answer of questions no q. how to take online courses I am from Nepal